OpenStudy (anonymous):

The coordinates of n ABC are A(0, 5), B(8, 20), and C(0, 26). Find the length of each side and the perimeter of n ABC. Then fi nd the perimeter of the triangle formed by connecting the three midsegments of n ABC.

OpenStudy (anonymous):

Get each side from L = sqrt( (x1-x2)^2 + (y1-y2))^2 So AB is L = sqrt((0-8)^2 + (5-20)^2) L = sqrt(64+ 225) =17 Do same for BC and CA Perimeter is the sum of the three sides.

OpenStudy (wolf1728):

cayytherine I don't know if you've already determined the answer but I have made a graphic of the triangle and determined the lengths of all 3 sides. If nothing else, see the attached graphic just to check your answers.

OpenStudy (wolf1728):

I can't believe I made a mistake. Line AC = 21 and NOT 25. So the perimeter = 48 (The mistake I made was simply subtraction - no trig no geometry) LOL